# Geological Everything Model 3D: A Promptable Foundation Model for Unified and Zero-Shot Subsurface Understanding

### v1

@@ Abstract
Understanding Earth’s subsurface is critical for energy transition, natural hazard mitigation, and planetary science. Yet subsurface analysis remains fragmented, with separate models required for structural interpretation, stratigraphic analysis, geobody segmentation, and property modeling—each tightly coupled to specific data distributions and task formulations. We introduce the Geological Everything Model 3D (GEM), a unified generative architecture that reformulates all these tasks as prompt-conditioned inference along latent structural frameworks derived from subsurface imaging. This formulation moves beyond task-specific models by enabling a shared inference mechanism, where GEM propagates human-provided prompts—such as well logs, masks, or structural sketches—along inferred structural frameworks to produce geologically coherent outputs. Through this mechanism, GEM achieves zero-shot generalization across tasks with heterogeneous prompt types, without retraining for new tasks or data sources. This capability emerges from a two-stage training process that combines self-supervised representation learning on large-scale field seismic data with adversarial fine-tuning using mixed prompts and labels across diverse subsurface tasks. GEM demonstrates broad applicability across surveys and tasks, including Martian radar stratigraphy analysis, structural interpretation in subduction zones, full seismic stratigraphic interpretation, geobody delineation, and property modeling. By bridging expert knowledge with generative reasoning in a structurally aware manner, GEM lays the foundation for scalable, human-in-the-loop geophysical AI—transitioning from fragmented pipelines to a vertically integrated, promptable reasoning system.
@@ 1 Introduction
Understanding Earth’s subsurface is foundational to some of society’s most pressing challenges: sustainable energy exploration [1, 2], geological carbon storage [3, 4], natural hazard prediction [5, 6], and even planetary stratigraphy on planets [7, 8]. In recent years, Artificial Intelligence (AI) has emerged as a a transformative paradigm for subsurface analysis [9, 10, 11, 12, 13]. However, current implementations remain fragmented, brittle, and narrowly scoped. Current workflows depend on task-specific models—each trained on narrow datasets and bound to a single survey region—to perform structural interpretation [14, 15, 16], stratigraphic analysis[17, 18, 19], geobody segmentation [20, 21, 22, 23] or property modeling [24, 25, 26]. This fragmentation hinders generalization and requires costly retraining for every new task or area. While foundation models have revolutionized vision and language[27, 28, 29, 30, 31], no such model has emerged for subsurface imaging. This is not simply a matter of scale, but of epistemology: geological imaging data is sparse, indirect, and structurally entangled—lacking the high-density semantic cues that foundation models in other fields rely on. Bridging this domain gap calls for rethinking how we represent, prompt, and reason about the Earth’s interior—not as a patchwork of disconnected labels, but as a unified structure-conditioned generative process.
@@
Several recent models have made progress toward more adaptable subsurface interpretation. The Seismic Foundation Model (SFM), based on masked autoencoders, achieves strong cross-survey results on horizon and fault detection, but remains restricted to dense supervision and narrow task formats [31, 40]. The Vision Foundation Model (VFM), adapted from the Segment Anything Model (SAM), enables prompt-guided segmentation using sparse masks or points, yet it is limited to 2D slices and requires task-specific fine-tuning for most applications [30, 39]. Other methods have incorporated domain priors—such as physics-informed loss functions [33, 34, 35], stratigraphic simulation constraints [36, 37, 19], and weak supervision strategies [41, 20]—but struggle to accommodate heterogeneous prompt types or task modalities within a shared architecture. Moreover, most models view interpretation as per-voxel classification or segmentation, lacking a structural representation of geology that can support consistent reasoning across scales and tasks. These limitations reveal the absence of a unified, promptable, and structure-aware modeling paradigm for subsurface interpretation.
@@
In response to these limitations, we introduce the Geological Everything Model 3D (GEM), a unified generative framework that reformulates subsurface interpretation as prompt-conditioned inference over latent structural representations. Instead of treating tasks like fault detection, geobody segmentation, or property modeling as separate problems requiring bespoke models, GEM abstracts them as different forms of dimensional completion guided by human-provided prompts—such as well logs, masks, or structural sketches. At its core, GEM learns to propagate these prompts through a structurally coherent latent space derived from the input volume, enabling geologically consistent outputs across diverse task types. This formulation unifies a wide range of interpretation goals within a single model, while maintaining flexibility for expert interaction and scalability across surveys. By shifting the modeling paradigm from task-specific classification to structure-aware generation, GEM offers a foundation for zero-shot, prompt-driven, human-in-the-loop geological reasoning.
@@
GEM is trained in two stages to enable broad generalization without task-specific retraining. First, a self-supervised pretraining stage leverages masked representation learning over more than 500 field seismic volumes to learn structure-sensitive features without requiring labels. Then, adversarial fine-tuning is performed using synthetic data with diverse prompt-label pairs across structural, stratigraphic, geobody, and physical property domains. This combination allows GEM to align prompts with latent geological frameworks, supporting flexible input types—including points, lines, masks, and well logs—and producing task-consistent outputs with minimal supervision. Once trained, GEM generalizes across surveys, data modalities, and interpretation goals, achieving zero-shot performance without architectural modification or additional fine-tuning. We evaluate GEM across a diverse set of interpretation tasks and geological settings, demonstrating broad applicability and strong zero-shot generalization. Without retraining, GEM performs high-resolution Martian radar stratigraphy analysis, interprets complex 3D fault systems in subduction zones, conducts full seismic stratigraphic interpretation, segments geobodies such as salt domes and channels, and models elastic properties across surveys with sparse well control. These applications span planetary exploration, structural geology, reservoir modeling, and seismic inversion—tasks that traditionally require disconnected models, domain-specific workflows, and extensive retraining at each stage. In contrast, GEM consolidates structural interpretation, stratigraphic analysis, geological modeling, and petrophysical prediction within a single generative framework. This unified approach eliminates the need for task-specific pipelines and mitigates error accumulation across stages, offering a more stable, scalable, and interpretable alternative. By integrating expert prompts with structure-aware reasoning, GEM bridges geoscientific insight and generative modeling, establishing a new paradigm for human-guided subsurface AI. This work makes three key contributions:
@@ 2.1 GEM: A Unified and Prompt-Driven Reasoning Framework for Subsurface imaging
GEM is designed as a unified, prompt-driven foundation model for diverse subsurface imaging tasks. It reformulates structural interpretation, stratigraphic analysis, geobody segmentation, and physical property modeling as conditional generative processes, driven by sparse user-provided prompts and executed over a shared latent geological framework. This formulation allows GEM to handle multiple interpretation goals within a single architecture. As illustrated in Figure 1a, GEM supports a broad spectrum of tasks—including fault and horizon delineation, relative geological time (RGT) estimation, geobody segmentation (e.g., channels, salt, volcanic units), and property modeling (e.g., impedance, gamma ray, lithology)—all through a consistent, prompt-conditioned inference mechanism.
@@
To enable this unified and interactive inference, GEM is trained by a two-stage strategy (Figure 1 (b, c)). During self-supervised pretraining, the model learns structure-sensitive representations from over 500 unlabeled field seismic volumes via masked modeling. In parallel, a perception network is trained to extract high-level geological features—such as faults, stratigraphy, and geobodies—which are later used as frozen feature extractors to provide structure-aware supervision during fine-tuning via perceptual loss. In the supervised fine-tuning stage, GEM is optimized using heterogeneous prompt-label pairs and trained with adversarial and perceptual losses. These losses encourage geological plausibility across outputs, while enabling generalization to unseen tasks, data modalities, and regions. For large-scale 3D deployment, GEM employs a purely convolutional architecture without self-attention layers (Figure 12), ensuring efficient inference without compromising spatial coherence. The training data preparation is detailed in Appendices B, and the model architecture and training process are described in the Methods section.
By integrating structure-conditioned learning with prompt-guided generation and expert interaction, GEM establishes a new modeling paradigm for subsurface interpretation. It transitions the traditional fragmented pipeline into a unified, vertically integrated reasoning system. The following sections demonstrate GEM’s zero-shot performance across structural, stratigraphic, geobody, and property interpretation tasks in both terrestrial and planetary settings.
@@ 2.2 Emergent Structure-Aware Reasoning in GEM
Figure 2 (a) illustrates this emergent structural understanding of subsurface during inference. To visualize GEM’s internal representations computed from input seismic volumes (a-1, a-3), we extract feature maps from the first upsampling stage and apply Principal Component Analysis (PCA) along the channel dimension to reduce them to three principal components. These reduced features are then projected into the RGB color space, resulting in visualizations (a-2, a-4) that consistently and coherently highlight critical geological elements such as faults, unconformities, and marker horizons. The spatial patterns of these features closely align with expert interpretations, as indicated by the dashed curves in (a-1, a-2). Note that the expert annotations are not used as inputs or prompts. This observation confirms that GEM can infer high-level geological concepts through its internal representations, without any explicit structural guidance during inference.
@@
To better understand how prompts guide internal reasoning, we freeze the pretrained GEM model and attach a lightweight multilayer perceptron (MLP) to probe its intermediate features during forward inference. As shown in Figure 2 (c), we examine how different prompt types (a 2D slice or a vertical well-log) trigger distinct patterns of feature activation. These MLP-projected visualizations reveal how prompt information propagates through successive network modules ( to ), gradually expanding from local prompt regions into structurally coherent representations. The slice prompt activates localized features that spread laterally along a geologic stratum, while the well-log prompt initiates vertical propagation that aligns with stratigraphic continuity. This behavior mirrors expert interpretation: starting from sparse observations, GEM progressively extrapolates geologically consistent structures that respect boundaries and relationships encoded in the subsurface image. Such prompt-aware, structure-conditioned reasoning underlies GEM’s ability to unify diverse interpretation tasks—ranging from structural delineation and geobody segmentation to property modeling—within a single, generalizable framework.
@@ 2.3 Structural and Stratigraphic Interpretation
Interpreting subsurface structures and stratigraphy remains fundamentally constrained by the highly heterogeneous and non-stationary nature of sedimentary systems. These complexities arise not only from primary depositional heterogeneities but are further compounded by post-depositional geological processes such as faulting, folding, karstification, magmatic intrusions, salt tectonics, and diagenesis. Such processes disrupt original stratigraphic architectures through dislocation, truncation, and deformation, leading to disordered and discontinuous sequences that deviate significantly from idealized layer-cake models. Subsurface imaging techniques, such as seismic imaging, are further limited by acquisition bandwidth, resolution, signal-to-noise ratio, and inaccuracies in velocity modeling—factors that reduce imaging fidelity, and semantic richness, ultimately contributing to significant interpretational ambiguity.
@@
GEM addresses this challenge by serving as a human-in-the-loop geophysical foundation model that enables expert-guided interpretation across diverse geological settings. It allows interpreters to interactively compute fault surfaces, horizons, or RGT volumes through multimodal prompts. We validate GEM across a diverse set of challenges, including décollement and thrust fault interpretation in the Hikurangi margin [42], bottom simulation reflector (BSR) tracking, Martian ice reflector delineation using SHARAD radargrams [43], unconformity surface extraction, and RGT volume estimation or fully stratigraphic interpretation in the Poseidon and Delft surveys [44]. Additional case studies are provided in the Appendices (Figure 9), demonstrating GEM’s versatility in both terrestrial and extraterrestrial seismic datasets.
@@ 2.3.1 Quantitative evaluation metric
Before presenting these case studies in detail, we first present GEM’s overall quantitative performance on both synthetic and field datasets. The synthetic dataset consists of 100 volumes, of size , generated using the same procedures as those used for training. For field data, the quantitative test set includes 20 expert-annotated slices from the Hikurangi Margin and Baiyun surveys (Figure 9 (b)).
@@ 2.3.2 Fault interpretation of the Hikurangi subduction zone
We focus on the task of identifying and characterizing specialized fault structures along the Hikurangi margin [42], a region that poses substantial challenges for automated interpretation. This margin is heavily influenced by the subduction of seamounts, which introduces complex, non-planar fault geometries and heterogeneous lithological conditions. These features significantly complicate fault interpretation, particularly for décollements, thrust systems, and fluid-rich sedimentary lenses. The intricate spatial relationships between structures, combined with subtle seismic expressions, often hinder reliable detection using conventional techniques.
Existing methods, such as FaultNet and FaultSeg+, have been widely adopted in industry practice. However, as illustrated in Appendices Figure 9 (a), these models struggle in this context—they tend to overfit to secondary normal faults and frequently miss the key structural elements that control slip behavior along the margin. Consequently, practitioners often resort to manual interpretation, which entails tracing faults across numerous 2D seismic sections followed by interpolation into 3D surfaces. This process can take weeks to complete [42], and it often lacks geometric precision and fails to capture fine-scale fault morphology critical for mechanistic analysis.
@@
These enhanced results provide crucial support for interpreting the mechanical framework of the Hikurangi margin. For instance, the data reveal that the Pāpaku fault—developing along the trailing edge of a subducting seamount—forms a fluid-rich sedimentary lens that maintains high pore pressure and impedes overlying plate collapse. Similarly, the spatial coincidence between an ancient sedimentary lens beneath the Tuaheni fault system and the 2014 slow slip zone suggests that such structures exert sustained control on slow slip activity. GEM facilitates the extraction and 3D reconstruction of these complex fault systems with high precision, enabling geoscientists to more efficiently and accurately delineate the fault geometries that underpin these observations. By reducing reliance on time-consuming manual interpretation, GEM accelerates the process of linking seismic structure to geomechanical behavior, thereby enhancing our ability to interpret the role of fault architecture and fluid conditions in controlling shallow slow slip events.

### v2

@@ Abstract
Subsurface understanding is critical for energy transition, natural hazard mitigation, and planetary science. Yet subsurface analysis remains fragmented, with separate models required for structural interpretation, stratigraphic analysis, geobody segmentation, and property modeling—each tightly coupled to specific data distributions and task formulations. We introduce the Geological Everything Model 3D (GEM), a unified generative architecture that reformulates all these tasks as prompt-conditioned inference along latent structural frameworks derived from subsurface imaging. This formulation moves beyond task-specific models by enabling a shared inference mechanism, where GEM propagates human-provided prompts—such as well logs, masks, or structural sketches—along inferred structural frameworks to produce geologically coherent outputs. Through this mechanism, GEM achieves zero-shot generalization across tasks with heterogeneous prompt types, without retraining for new tasks or data sources. This capability emerges from a two-stage training process that combines self-supervised representation learning on large-scale field seismic data with adversarial fine-tuning using mixed prompts and labels across diverse subsurface tasks. GEM demonstrates broad applicability across surveys and tasks, including Martian radar stratigraphy analysis, structural interpretation in subduction zones, full seismic stratigraphic interpretation, geobody segmentation, and property modeling. By bridging expert knowledge with generative reasoning in a structurally aware manner, GEM lays the foundation for scalable, human-in-the-loop geophysical AI—transitioning from fragmented pipelines to a vertically integrated, promptable reasoning system.
@@ 1 Introduction
Subsurface understanding is foundational to some of society’s most pressing challenges: sustainable energy exploration [1, 2], geological carbon storage [3, 4], natural hazard prediction [5, 6] and stratigraphy [7, 8]. In recent years, Artificial Intelligence (AI) has emerged as a transformative paradigm for subsurface analysis [9, 10, 11, 12, 13]. However, current implementations remain fragmented, brittle, and narrowly scoped. Most workflows depend on task-specific models—each trained on narrow datasets and bound to a single survey region—to perform structural interpretation [14, 15, 16], stratigraphic analysis[17, 18, 19], geobody segmentation [20, 21, 22, 23] or property modeling [24, 25, 26, 27]. This fragmentation hinders generalization and requires costly retraining for every new task or area. While foundation models have revolutionized vision and language[28, 29, 30, 31, 32], no such model has emerged for subsurface imaging. This is not simply a matter of scale, but of epistemology: geological imaging data is sparse, indirect, and structurally entangled—lacking the high-density semantic cues that foundation models in other fields rely on. Bridging this domain gap calls for rethinking how we represent, prompt, and reason about the planetary interior—not as a patchwork of disconnected labels, but as a unified structure-conditioned generative process.
@@
Several recent models have made progress toward more adaptable subsurface interpretation [41, 42]. The Seismic Foundation Model (SFM), based on masked autoencoders, achieves strong cross-survey results on horizon and fault detection, but remains restricted to dense supervision and narrow task formats [32, 43]. The Vision Foundation Model (VFM), adapted from the Segment Anything Model (SAM), enables prompt-guided segmentation using sparse masks or points, yet it is limited to 2D slices and requires task-specific fine-tuning for most applications [31, 40]. Other methods have incorporated domain priors—such as physics-informed loss functions [34, 35, 36], stratigraphic simulation constraints [37, 38, 19], and weak supervision strategies [44, 20]—but struggle to accommodate heterogeneous prompt types or task modalities within a shared architecture. Moreover, most models view interpretation as per-voxel classification or segmentation, lacking a structural representation of geology that can support consistent reasoning across scales and tasks. These limitations reveal the absence of a unified, promptable, and structure-aware modeling paradigm for subsurface interpretation.
@@
In response to these limitations, we introduce the Geological Everything Model 3D (GEM), a unified generative framework that reformulates subsurface interpretation as prompt-conditioned inference over latent structural representations. Instead of treating tasks like structural interpretation, geobody segmentation, or property modeling as separate problems requiring bespoke models, GEM abstracts them as different forms of dimensional completion guided by human-provided prompts—such as well logs, masks, or structural sketches. At its core, GEM learns to propagate these prompts through a structurally coherent latent space derived from the input subsurface volume, enabling geologically consistent outputs across diverse task types. This formulation unifies a wide range of interpretation goals within a single model, while maintaining flexibility for expert interaction and scalability across surveys. By shifting the modeling paradigm from task-specific classification to structure-aware generation, GEM offers a foundation for zero-shot, prompt-driven, human-in-the-loop geological reasoning.
@@
GEM is trained in two stages to enable broad generalization without task-specific retraining. First, a self-supervised pretraining stage leverages masked representation learning over more than 500 field seismic volumes to learn structure-sensitive features without requiring labels. Then, adversarial fine-tuning is performed using synthetic data with diverse prompt-label pairs across structural, stratigraphic, geobody, and physical property domains. This combination allows GEM to align prompts with latent geological frameworks, supporting flexible input types—including lines, masks, and well logs—and producing task-consistent outputs with minimal supervision. Once trained, GEM generalizes across surveys, data modalities, and interpretation goals, achieving zero-shot performance without architectural modification or additional fine-tuning. We evaluate GEM across a diverse set of tasks and geological settings, demonstrating broad applicability and strong zero-shot generalization. Without retraining, GEM performs high-resolution Martian radar stratigraphy analysis, interprets complex 3D fault systems in subduction zones, conducts full seismic stratigraphic interpretation, segments geobodies such as salt domes and channels, and models properties across surveys with sparse well control. These applications span planetary exploration, structural geology and property modeling—tasks that traditionally require disconnected models, domain-specific workflows, and extensive retraining at each stage. In contrast, GEM consolidates structural interpretation, stratigraphic analysis, geological modeling, and property prediction within a single generative framework. This unified approach eliminates the need for task-specific pipelines and mitigates error accumulation across stages, offering a more stable, scalable, and interpretable alternative. By integrating expert prompts with structure-aware reasoning, GEM bridges geoscientific insight and generative modeling, establishing a new paradigm for human-guided subsurface AI. This work makes three key contributions:
@@ 2.1 GEM: A Unified and Prompt-Driven Reasoning Framework for Subsurface imaging
GEM is designed as a unified, prompt-driven foundation model for diverse subsurface imaging tasks. It reformulates structural interpretation, stratigraphic analysis, geobody segmentation, and physical property modeling as conditional generative processes, driven by sparse user-provided prompts and executed over a shared latent geological framework. This formulation allows GEM to handle multiple interpretation goals within a single architecture. As illustrated in Figure 1a, GEM supports a broad spectrum of tasks—including fault and horizon extraction, relative geological time (RGT) estimation, geobody segmentation (e.g., channels, salt, volcanic units), and property modeling (e.g., impedance, gamma ray, lithology)—all through a consistent, prompt-conditioned inference mechanism.
@@
To enable this unified and interactive inference, GEM is trained by a two-stage strategy (Figure 1 (b, c)). During self-supervised pretraining, the model learns structure-sensitive representations from over 500 unlabeled field seismic volumes via masked modeling. In parallel, a perception network is trained to extract high-level geological features—such as faults, stratigraphy, and geobodies—which are later used as frozen feature extractors to provide structure-aware supervision during fine-tuning via perceptual loss. In the supervised fine-tuning stage, GEM is optimized using heterogeneous prompt-label pairs and trained with adversarial and perceptual losses. These losses encourage geological plausibility across outputs, while enabling generalization to unseen tasks, data modalities, and regions. For large-scale 3D deployment, GEM employs a purely convolutional architecture (Figure 12), ensuring efficient inference without compromising spatial coherence. The training data preparation is detailed in Appendices B, and the model architecture and training process are described in the Methods section.
@@ 2.2 Emergent Structure-Aware Reasoning in GEM
Figure 2 (a) illustrates this emergent structural understanding of subsurface during inference. To visualize GEM’s internal representations computed from input seismic volumes (a-1, a-3), we extract feature maps from the first upsampling stage and apply Principal Component Analysis (PCA) along the channel dimension to reduce them to three principal components. These reduced features are then projected into the RGB color space, resulting in visualizations (a-2, a-4) that consistently and coherently highlight critical geological elements such as faults, unconformities, and marker horizons. The spatial patterns of these features closely align with expert interpretations, as indicated by the dashed curves in (a-1, a-3). Note that the expert annotations are not used as inputs or prompts. This observation confirms that GEM can infer high-level geological concepts through its internal representations, without any explicit structural guidance during inference.
@@
To better understand how prompts guide internal reasoning, we freeze the pretrained GEM model and attach a lightweight multilayer perceptron (MLP) to probe its intermediate features during forward inference. As shown in Figure 2 (c), we examine how different prompt types (a 2D mask or a vertical well-log) trigger distinct patterns of feature activation. These MLP-projected visualizations reveal how prompt information propagates through successive network modules ( to ), gradually expanding from local prompt regions into structurally coherent representations. The mask prompt activates localized features that spread laterally along a geologic stratum, while the well-log prompt initiates vertical propagation that aligns with stratigraphic continuity. This behavior mirrors expert interpretation: starting from sparse observations, GEM progressively extrapolates geologically consistent structures that respect boundaries and relationships encoded in the subsurface image. Such prompt-aware, structure-conditioned reasoning underlies GEM’s ability to unify diverse interpretation tasks—ranging from structural delineation and geobody segmentation to property modeling—within a single, generalizable framework.
@@ 2.3 Structural and Stratigraphic Interpretation
Interpreting subsurface structures and stratigraphy remains fundamentally constrained by the highly heterogeneous and non-stationary nature of sedimentary systems. These complexities arise not only from primary depositional heterogeneities but are further compounded by post-depositional geological processes such as faulting, folding, karstification, magmatic intrusions, salt tectonics, and diagenesis. Such processes disrupt original stratigraphic architectures through deformation dislocation and truncation leading to disordered and discontinuous sequences that deviate significantly from idealized layer-cake models. Subsurface imaging techniques, such as seismic imaging, are further limited by acquisition bandwidth, resolution, signal-to-noise ratio, and inaccuracies in velocity modeling—factors that reduce imaging fidelity, and semantic richness, ultimately contributing to significant interpretational ambiguity.
@@
GEM addresses this challenge by serving as a human-in-the-loop geophysical foundation model that enables expert-guided interpretation across diverse geological settings. It allows interpreters to interactively compute fault surfaces, horizons, or RGT volumes through multimodal prompts. We validate GEM across a diverse set of challenges, including décollement and thrust fault interpretation in the Hikurangi margin [45], bottom simulation reflector (BSR) tracking, Martian ice reflector extraction using SHARAD radargrams [46], unconformity surface extraction, and RGT volume estimation or fully stratigraphic interpretation in the Poseidon and Delft surveys [47]. Additional case studies are provided in the Appendices (Figure 9), demonstrating GEM’s versatility in both terrestrial and extraterrestrial datasets.
@@ 2.3.1 Quantitative evaluation metric
Before presenting these case studies in detail, we first present GEM’s overall quantitative performance on both synthetic and field datasets. The synthetic dataset consists of 100 volumes, of size . For field dataset includes 20 expert-annotated slices from the Hikurangi Margin and Baiyun surveys (Figure 9 (b)).
@@ 2.3.2 Fault interpretation of the Hikurangi subduction zone
We focus on the task of identifying and characterizing specialized, complex fault structures along the Hikurangi margin[45], a region that poses substantial challenges for automated interpretation. This subduction zone margin is heavily influenced by collisions with rough basement associated with seamounts on the subduction plates. The relief from subducting seamounts introduces complex, non-planar fault geometries, heterogeneous lithologies created by the influence of seamounts on basin deposition, and heterogeneous sediment compaction associated with stress shadows created by seamounts. These features significantly complicate fault interpretation, particularly for décollements, thrust systems, and fluid-rich sedimentary lenses associated with seamount induced stress shadows. The intricate spatial relationships between structures, combined with subtle seismic expressions, often hinder reliable detection using more conventional automated techniques.
Existing methods, such as FaultNet and FaultSeg+, have been widely adopted in industry practice. However, as illustrated in Appendices Figure 9 (a), these models struggle in this challenging subduction zone setting—they tend to overfit to secondary normal faults and frequently miss the key structural elements that control slip behavior along the margin. Consequently, practitioners often resort to manual interpretation, which entails tracing faults across numerous 2D seismic sections followed by interpolation into 3D surfaces. This process can take weeks to complete[45], and it often lacks geometric precision and fails to capture fine-scale fault morphology critical for mechanistic analysis.
@@
These enhanced results provide crucial support for interpreting the mechanical framework of the Hikurangi margin. For instance, the data reveals that the Pāpaku fault—developing along the trailing edge of a subducting seamount—forms a fluid-rich sedimentary lens in the stress shadow of the seamount that maintains high pore pressure and impedes overlying plate collapse. Similarly, the spatial coincidence between ancient larger, earlier formed sedimentary lens beneath the Tuaheni fault system and the 2014 slow slip zone suggests that such under-compacted sediment lenses exert sustained control on slow slip activity. GEM facilitates the extraction and 3D reconstruction of these complex fault systems with high precision, enabling geoscientists to more efficiently and accurately delineate the fault geometries that underpin these observations. By reducing reliance on time-consuming manual interpretation, GEM accelerates the process of linking seismic structure to geomechanical behavior, thereby enhancing our ability to interpret the role of fault architecture and fluid conditions in controlling shallow slow slip events.